# Strategy for identifying Vera C. Rubin Observatory kilonova candidates for targeted gravitational-wave searches

### v1

@@ 4 Identification of kilonova candidates in Rubin with Fink
In this work, we aim to create a search strategy for kilonova candidates directly from Rubin data communicated to brokers. In Section  4.1, we introduce Fink which is one of the Rubin Community brokers, that aims to enable a broad range of science cases (A. Möller et al., 2021). In Section 4.2, we translate the simulations introduced in Section 3.2 into the data that the brokers will receive and study the kilonova candidate properties. We then devise different selection criteria for kilonova candidates in Rubin and test them using archival data from the Zwicky Transient Facility (ZTF) in Section 4.3.
@@ 4.1 Fink
The Fink community broker is designed to process large-scale time-domain alert streams, filtering, aggregating, enriching, and redistributing data in real time (A. Möller et al., 2021). Built as part of the Rubin Observatory Community Broker ecosystem, Fink was conceived to handle the full alert stream of the Vera C. Rubin Observatory LSST (Ž. Ivezić et al., 2019). Its modular architecture allows independent science teams to develop filtering and classification routines (“science modules”) for specific astrophysical phenomena. Once integrated, these modules operate within the centralized broker infrastructure, making enriched alert data accessible through its web portal [^6] https://fink-portal.org/ and data transfer services [^7] https://fink-portal.org/download.
Since late 2019, Fink has been operating on the ZTF public stream (E. C. Bellm et al., 2018), serving as a precursor to LSST operations. ZTF alerts that pass basic quality criteria are processed through science modules, which add value through catalog cross-matching, statistical analyses, and machine-learning classification routines [^8] https://fink-broker.readthedocs.io/en/latest/broker/science_modules/. The breadth of implemented modules reflects the diversity of the Fink community, spanning supernovae (M. Leoni et al., 2022; A. Möller & T. de Boissière, 2020; A. Möller et al., 2025), kilonovae (B. Biswas et al., 2023), tidal disruption events (M. Llamas Lanza et al., 2025), gamma-ray bursts (M. Masson & J. Bregeon, 2024), microlensing (M. Ban et al., 2025), solar system objects (R. Le Montagner et al., 2023), hostless transients (P. J. Pessi et al., 2024), and multi-class classifiers (B. M. O. Fraga et al., 2024), among others. Each module is developed independently, but their outputs are publicly available, with transparent assumptions to enable informed community use.
@@ 4.2.1 Rubin alert stream
Rubin LSST will produce an alert stream. The alert stream will contain detections from the Difference Image Analysis pipeline with signal-to-noise (SNR) greater than 5 (E. Bellm et al., 2020). For a given object, defined by an event in the same coordinates, if two alerts (SNR>5>5) are emitted in different nights, the latest alert will also contain the history of forced photometry [^9] https://community.lsst.org/t/alert-information-contained-with-the-two-or-three-visits-per-night/10068/ and https://ldm-612.lsst.io/. We show examples of a kilonova light-curve as seen by the broker, and compare to the full detected Rubin light-curve in Figure 1.
@@ 4.2.2 Rubin kilonova properties
We find that the two most-common band-passes where kilonovae are detected in Rubin are $i$ and $r$ followed, with half of detections by the filter $z$, and then with small percentages in the order $y,g,u$. This is surprising as previous works showed that $g$ and $i$ were the most probable filters (I. Andreoni et al., 2022a) for kilonova detection. We attribute this difference to the use of updated observing strategies for Rubin (version 4.3 and Ocean) in this work.
The Rubin Science Verification strategy, shaped through a strong community effort [^10] https://docs.google.com/document/d/1sssH6buVEfXj7lEOnckuXC4sg-JxNlSlgcKnRffb0fk/edit?usp=sharing, has been recently refined [^11] https://sitcomtn-005.lsst.io to prioritize template acquisition in the $i$ band, particularly in regions of the sky where the LIGO-Virgo-KAGRA network has maximum antenna sensitivity (cf. H.-Y. Chen et al., 2017). Consequently, the $i$ band is expected to play a central role in the identification of kilonova candidates.
@@ 4.2.3 Time between explosion and first detection
Figure 4 shows the distribution of times between when a kilonova candidate is first detected by Fink/Rubin and a) the time of maximum light and b) the time when the neutron star merger occurred (‘explosion time’). We find that most kilonovae are detected around 1.5 days after maximum light and between 1-2 days after explosion. This figure provides us with an estimate of the typical length of data that would have to be searched in order to be confident that the searched data includes the time of the gravitational-wave emission. This is significantly longer than the on-source windows of a few seconds that are used in targeted gravitational-wave searches for gamma-ray bursts (R. Abbott et al., 2021b, 2022), and is more similar to targeted searches for core-collapse supernovae (B. P. Abbott et al., 2020c; M. J. Szczepańczyk et al., 2024; A. G. Abac et al., 2025e). We discuss implications for targeted gravitational-wave searches further in Section 5.
@@ 4.3 Fink filter for Rubin kilonova candidates
In the following, we propose different Fink broker filters to select kilonova candidates. As a guideline, we take the filters created by the Fink and GRANDMA Collaborations (S. Antier et al., 2020). Since 2021, they have developed 3 different kilonova candidate filters based on: (i) Machine Learning algorithm scores (B. Biswas et al., 2023), (ii) Potential host-galaxy cross-match using Mangrove catalogue (J.-G. Ducoin et al., 2020) and (iii) decline rate of the light-curve (I. Andreoni et al., 2021). There are also additional criteria to reject asteroids, known variables and transients. These filters have been deployed in Fink and have identified potential candidates (V. Aivazyan et al., 2022; I. Tosta e Melo et al., 2024).
We also take into account the criteria presented in D. Scolnic et al. (2018) and C. N. Setzer et al. (2019). The criteria are (1) have two alerts separated by at least 30 minutes to reject asteroids, (2) be observed in at least two filters with ${\rm SNR}\geq 5$ to characterise them, (3) include ${\rm SNR}\geq 5$ observations separated by no more than 25 days to reject long-lived events, (4) have at least one observation of the same location within 20 days prior to the first ${\rm SNR}\geq 5$ detection, and (5) have at least one observation of the same location within 20 days after the last ${\rm SNR}\geq 5$ detection.
@@
    Remove all sources cross-matched within 2” to Variable star catalogues such as the General Catalog of Variable Stars (GCVS; N. N. Samus’ et al., 2017) and the Variable Star Index (VSX; P. N. Kholopov et al., 1992).
@@
These are broad cuts that will select new detected objects, as well as avoid catalogued asteroids, variables and transients. However, a large fraction of the remaining alerts may be uncatalogued cataclysmic variables (CVs), supernovae, novae, Fast Blue Optical Transients (FBOTs) and gamma-ray burst afterglows (T. Barna et al., 2025).
We refine these cuts into two main selection families: all sky searches and galaxy-targeted searches as in I. Andreoni et al. (2020).
@@
  • All sky silver: Less than 3 days of detections and discarding any potential satellite or space debris using a tracklet (S. Karpov & J. Peloton, 2023).
@@ 4.3.1 Test with ZTF alert stream
To validate our candidate selection methodology, we applied our filtering criteria to alerts processed by the Fink broker (A. Möller et al., 2021) from the Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2018). ZTF surveys the northern hemisphere with a limiting magnitude of $\approx 21$ and generates alerts for detections $g,r$ band-passes, which are subsequently processed by Fink.
@@
The Galaxy-targeted criteria substantially reduce candidate numbers by restricting selection to transients in proximity to local volume galaxies. However, incomplete galaxy catalogs may introduce selection biases. Our current analysis utilizes the Mangrove catalog, which is limited to 200 Mpc. The ongoing 4MOST 4-metre Multi-Object Spectroscopic Telescope Hemisphere Survey (E. N. Taylor et al., 2023) will provide more complete galaxy samples with spectroscopic redshifts for $z<0.1$.
@@
Both All-Sky Gold/Silver and Galaxy Gold criteria restrict candidates to fewer than 35 detections per month for ZTF alert volumes and depth. For Rubin LSST this may represent a lower threshold. Together with the filtering, in particular at the beginning of the survey, it will be crucial to perform visual inspection of the candidates, in particular the thumbnails which can show previous sources pointing towards variable stars, events located in the centre of galaxies pointing towards AGNs, as well as other artefacts from image subtraction.
@@ 4.4 Extending the kilonova search with Rubin proprietary data
These findings supports the adoption of a lower SNR threshold for kilonova searches in the Rubin ToO program (I. Andreoni et al., 2024). However, for blind searches a compromise between a higher threshold to eliminate spurious detections and the number of detections per light-curve should be considered.
@@ 4.5 Optical follow-up scenarios
Photometric follow-up to enhance sampling of kilonova light curve evolution represents an optimal strategy. However, the peak magnitudes of the kilonovae in our volume, shown in Figure 2, reveal that many simulated events are quite faint. Current wide-field surveys such as ZTF and LS4 (E. C. Bellm et al., 2018; A. A. Miller et al., 2025) achieve single-exposure limiting magnitudes of approximately 21.5 mag. Our simulations indicate that only 1–2% of kilonovae would reach peak brightnesses within these detection limits. Furthermore, given the limited sky overlap between ZTF (in the Northern hemisphere) and the Rubin Observatory (in the Southern hemisphere), this already small fraction would be substantially reduced in practice potentially by a factor larger than 3. Therefore, the single-exposure detection thresholds of the current ZTF alert stream limit its ability to provide substantial additional photometric coverage for the faint kilonova population accessible to Rubin.
Spectroscopic confirmation of kilonova candidates will prove essential for distinguishing between kilonovae and other variables and transients. However, given the expected peak magnitudes of kilonova candidates in Rubin, there are substantial observational challenges. We find that 2m telescopes reaching 191919.5mag19.5~mag (such as the ANU 2.3m telescope equipped with WiFeS; M. Dopita et al., 2007) would be able to follow less than $0.05\%$ of the candidates. A more pragmatic approach involves targeting the host galaxy of each candidate to determine its redshift, thereby enabling estimation of the absolute magnitude and intrinsic luminosity of the kilonova. This strategy provides crucial context for assessing the physical properties of potential kilonova events.
@@ 5 Targeted search for gravitational waves associated with FINK/Rubin kilonova candidates
The kilonova candidates identified in Section 4 provide astrophysical triggers for a targeted search for associated gravitational waves. These triggers provide an approximate time of the event, along with a sky location and (if using events with a known host galaxy) distance. In a targeted search for associated gravitational waves, this reduces the number of parameters that need to be searched over by up to four, allowing for a more sensitive search to be conducted (e.g., L. Z. Kelley et al., 2013; A. R. Williamson et al., 2014). This can ultimately lead to additional multi-messenger detections.
@@
There are several differences between a search for gravitational waves targeting kilonova candidates and one targeting gamma-ray bursts. Gamma-ray bursts associated with binary neutron star mergers are expected to occur at almost the exact time of the merger (cf. B. P. Abbott et al., 2017c, b), providing tight constraints on the time period of data to be searched (of order seconds) for gravitational waves (e.g., R. Abbott et al., 2021b, 2022). However, for a kilonova candidate detected by Rubin/Fink, the first detection will likely not correspond to the time of the associated neutron star merger. In the worst case scenario, a detection after peak brightness when the transient has already begun to fade, we know the peak must have been some time between the last non-detection and the first detection. This provides a window of order the cadence of the survey, which may be up to 5 days. In Section 4.2.3 we showed that most kilonova candidates identified by Fink will typically be detected one to two days after the neutron star merger. Targeted searches for gravitational waves associated with kilonovae will therefore typically need to search days of data, compared to seconds for searches targeting gamma-ray bursts. This leads to a situation that is more similar to targeted searches for gravitational waves from optically detected supernovae, which also typically have on-source windows of several days (e.g., B. P. Abbott et al., 2020c; M. J. Szczepańczyk et al., 2024; A. G. Abac et al., 2025e). This increases the computational cost of a targeted search, and it is likely that for many events detector data may not exist for the duration that needs to be searched due to the typical duty cycle of $\sim 70$% for gravitational-wave detectors (e.g., A. G. Abac et al., 2025f). We consider the case of GW170817/AT2017gfo as the best case scenario, that is, a kilonova detected before peak brightness, with a well sampled lightcurve, allowing for models to be fit to estimate the time of the merger. AT2017gfo was first detected in the optical approximately 12 hours after the merger time. From the optical data on AT2017gfo alone, the merger time can be estimated approximately (e.g., V. A. Villar et al., 2017).
@@
The other main difference between a gravitational-wave search targeting kilonova candidates and targeted searches for gravitational waves associated with gamma-ray bursts is that our list of targets will suffer from significant contamination. The majority of kilonova candidates identified by our selection criteria are not actually kilonovae associated with binary neutron star mergers, but rather other variables or transients such as supernovae and asteroids, with similar colors and magnitudes to kilonovae in one or two detections. The situation of a targeted search with significant contamination is similar to the search for gravitational waves associated with fast radio bursts, where some significant fraction (up to 100%) of fast radio bursts may originate from events that are not binary neutron star mergers (R. Abbott et al., 2023c).

### v2

@@ 4 Identification of kilonova candidates in Rubin with Fink
In this work, we aim to create a search strategy for kilonova candidates directly from Rubin data communicated to brokers. In Section  4.1, we introduce Fink which is one of the Rubin Community brokers, that aims to enable a broad range of science cases (Moller:2021MNRASFINK). In Section 4.2, we translate the simulations introduced in Section 3.2 into the data that the brokers will receive and study the kilonova candidate properties. We then devise different selection criteria for kilonova candidates in Rubin and test them using archival data from the Zwicky Transient Facility (ZTF) in Section 4.3.
@@ 4.1 Fink
The Fink community broker is designed to process large-scale time-domain alert streams, filtering, aggregating, enriching, and redistributing data in real time (Moller:2021MNRASFINK). Built as part of the Rubin Observatory Community Broker ecosystem, Fink was conceived to handle the full alert stream of the Vera C. Rubin Observatory LSST (Ivezic:2019ApJ). Its modular architecture allows independent science teams to develop filtering and classification routines (“science modules”) for specific astrophysical phenomena. Once integrated, these modules operate within the centralized broker infrastructure, making enriched alert data accessible through its web portal [^6] https://fink-portal.org/ and data transfer services [^7] https://fink-portal.org/download.
Since late 2019, Fink has been operating on the ZTF public stream (Bellm_2018), serving as a precursor to LSST operations. ZTF alerts that pass basic quality criteria are processed through science modules, which add value through catalog cross-matching, statistical analyses, and machine-learning classification routines [^8] https://fink-broker.readthedocs.io/en/latest/broker/science_modules/. The breadth of implemented modules reflects the diversity of the Fink community, spanning supernovae (Leoni:2022; Moller:2020; Moller:2025), kilonovae (Biswas:2023A&AFinkKilonova), tidal disruption events (LlamasLanza:2025), gamma-ray bursts (Masson:2024), microlensing (Ban:2025), solar system objects (Lemontagner:2023), hostless transients (Pessi:2024), and multi-class classifiers (Fraga:2024), among others. Each module is developed independently, but their outputs are publicly available, with transparent assumptions to enable informed community use.
@@ 4.2.1 Rubin alert stream
Rubin LSST will produce an alert stream. The alert stream will contain detections from the Difference Image Analysis pipeline with signal-to-noise (SNR) greater than 5 (ldm612). For a given object, defined by an event in the same coordinates, if two alerts (SNR>5>5) are emitted in different nights, the latest alert will also contain the history of forced photometry [^9] https://community.lsst.org/t/alert-information-contained-with-the-two-or-three-visits-per-night/10068/ and https://ldm-612.lsst.io/. We show examples of a kilonova light-curve as seen by the broker, and compare to the full detected Rubin light-curve in Figure 2.
@@ 4.2.2 Rubin kilonova properties
We find that the two most-common band-passes where kilonovae are detected in Rubin are $i$ and $r$ followed, with half of detections by the filter $z$, and then with small percentages in the order $y,g,u$. This is surprising as previous works showed that $g$ and $i$ were the most probable filters (Andreoni:2022ApJS) for kilonova detection. We attribute this difference to the use of updated observing strategies for Rubin (version 4.3 and Ocean) in this work.
The Rubin Science Verification strategy, shaped through a strong community effort [^10] https://docs.google.com/document/d/1sssH6buVEfXj7lEOnckuXC4sg-JxNlSlgcKnRffb0fk/edit?usp=sharing, has been recently refined [^11] https://sitcomtn-005.lsst.io to prioritize template acquisition in the $i$ band, particularly in regions of the sky where the LIGO-Virgo-KAGRA network has maximum antenna sensitivity (cf. Chen:2017ApJ). Consequently, the $i$ band is expected to play a central role in the identification of kilonova candidates.
@@ 4.2.3 Time between explosion and first detection
Figure 5 shows the distribution of times between when a kilonova candidate is first detected by Fink/Rubin and a) the time of maximum light and b) the time when the neutron star merger occurred (‘explosion time’). We find that most kilonovae are detected around 1.5 days after maximum light and between 1-2 days after explosion. This figure provides us with an estimate of the typical length of data that would have to be searched in order to be confident that the searched data includes the time of the gravitational-wave emission. This is significantly longer than the on-source windows of a few seconds that are used in targeted gravitational-wave searches for gamma-ray bursts (Abbott:2021ApJLVKGRBO3a; Abbott:2022ApJLVKGRBO3b), and is more similar to targeted searches for core-collapse supernovae (Abbott:2020PRDSNO1O2; Marek:2024PRDSNO3; Abac:2025ApJSN2023ixf). We discuss implications for targeted gravitational-wave searches further in Section 5.
@@ 4.3 Fink filter for Rubin kilonova candidates
In the following, we propose different Fink broker filters to select kilonova candidates. As a guideline, we take the filters created by the Fink and GRANDMA Collaborations (Antier:2020MNRAS). Since 2021, they have developed 3 different kilonova candidate filters based on: (i) Machine Learning algorithm scores (Biswas:2023A&AFinkKilonova), (ii) Potential host-galaxy cross-match using Mangrove catalogue (Ducoin:2020Mangrove) and (iii) decline rate of the light-curve (Andreoni:2021). There are also additional criteria to reject asteroids, known variables and transients. These filters have been deployed in Fink and have identified potential candidates (Aivazyan:2022; Melo:2024).
We also take into account the criteria presented in Scolnic:2018 and Setzer:2019MNRAS. The criteria are (1) have two alerts separated by at least 30 minutes to reject asteroids, (2) be observed in at least two filters with ${\rm SNR}\geq 5$ to characterise them, (3) include ${\rm SNR}\geq 5$ observations separated by no more than 25 days to reject long-lived events, (4) have at least one observation of the same location within 20 days prior to the first ${\rm SNR}\geq 5$ detection, and (5) have at least one observation of the same location within 20 days after the last ${\rm SNR}\geq 5$ detection.
@@
    Remove all sources cross-matched within 2” to Variable star catalogues such as the General Catalog of Variable Stars (GCVS; Samus:2017) and the Variable Star Index (VSX; Kholopov:1992).
@@
These are broad cuts that will select new detected objects, as well as avoid catalogued asteroids, variables and transients (including supernovae). However, a large fraction of the remaining alerts may be uncatalogued cataclysmic variables (CVs), supernovae, novae, Fast Blue Optical Transients (FBOTs) and gamma-ray burst afterglows (Barna:2025arXiv).
We refine these cuts into two main selection families: all sky searches and galaxy-targeted searches as in Andreoni:2020.
@@
  • All sky silver: Less than 3 days of detections and discarding any potential satellite or space debris using a tracklet (Karpov:2023).
@@ 4.3.1 Test with ZTF alert stream
To validate our candidate selection methodology, we applied our filtering criteria to alerts processed by the Fink broker (Moller:2021MNRASFINK) from the Zwicky Transient Facility (ZTF; Bellm_2018). ZTF surveys the northern hemisphere with a limiting magnitude of $\approx 21$ and generates alerts for detections $g,r$ band-passes, which are subsequently processed by Fink.
@@
The Galaxy-targeted criteria substantially reduce candidate numbers by restricting selection to transients in proximity to local volume galaxies. However, incomplete galaxy catalogs may introduce selection biases. Our current analysis utilizes the Mangrove catalog, which is limited to 200 Mpc. The ongoing 4MOST 4-metre Multi-Object Spectroscopic Telescope Hemisphere Survey (Taylor:2023) will provide more complete galaxy samples with spectroscopic redshifts for $z<0.1$.
@@
Figure 6: ZTF candidates that pass the different cut stages in the Fink broker. We show light-curves for these candidates in the first three columns and science cutouts (no subtraction) in the fourth column. Each row represents one of the sample cuts proposed in Section 4.3.
Both All-Sky Gold/Silver and Galaxy Gold criteria restrict candidates to fewer than 35 detections per month for ZTF alert volumes and depth. For Rubin LSST this may represent a lower threshold. Together with the filtering, in particular at the beginning of the survey, it will be crucial to perform visual inspection of the candidates, in particular the thumbnails which can show previous sources pointing towards variable stars, events located in the centre of galaxies pointing towards AGNs, as well as other artefacts from image subtraction. We show some examples of candidates that pass our various sample cuts in Figure 6.
@@ 4.4 Extending the kilonova search with Rubin proprietary data
These findings supports the adoption of a lower SNR threshold for kilonova searches in the Rubin ToO program (Andreoni:2024ToO). However, for blind searches a compromise between a higher threshold to eliminate spurious detections and the number of detections per light-curve should be considered.
@@ 4.5 Optical follow-up scenarios
Photometric follow-up to enhance sampling of kilonova light curve evolution represents an optimal strategy. However, the peak magnitudes of the kilonovae in our volume, shown in Figure 3, reveal that many simulated events are quite faint. Current wide-field surveys such as ZTF and LS4 (Bellm_2018; Miller:2025) achieve single-exposure limiting magnitudes of approximately 21.5 mag. Our simulations indicate that only 1–2% of kilonovae would reach peak brightnesses within these detection limits. Furthermore, given the limited sky overlap between ZTF (in the Northern hemisphere) and the Rubin Observatory (in the Southern hemisphere), this already small fraction would be substantially reduced in practice potentially by a factor larger than 3. Therefore, the single-exposure detection thresholds of the current ZTF alert stream limit its ability to provide substantial additional photometric coverage for the faint kilonova population accessible to Rubin.
Spectroscopic confirmation of kilonova candidates will prove essential for distinguishing between kilonovae and other variables and transients. However, given the expected peak magnitudes of kilonova candidates in Rubin, there are substantial observational challenges. We find that 2m telescopes reaching 191919.5mag19.5~mag (such as the ANU 2.3m telescope equipped with WiFeS; Dopita:2007) would be able to follow less than $0.05\%$ of the candidates. A more pragmatic approach involves targeting the host galaxy of each candidate to determine its redshift, thereby enabling estimation of the absolute magnitude and intrinsic luminosity of the kilonova. This strategy provides crucial context for assessing the physical properties of potential kilonova events.
@@ 5 Targeted search for gravitational waves associated with FINK/Rubin kilonova candidates
The kilonova candidates identified in Section 4 provide astrophysical triggers for a targeted search for associated gravitational waves. These triggers provide an approximate time of the event, along with a sky location and (if using events with a known host galaxy) distance. In a targeted search for associated gravitational waves, this reduces the number of parameters that need to be searched over by up to four, allowing for a more sensitive search to be conducted (e.g., Kelley:2013PRD; Williamson:2014PRDPyGRB). This can ultimately lead to additional multi-messenger detections.
@@
There are several differences between a search for gravitational waves targeting kilonova candidates and one targeting gamma-ray bursts. Gamma-ray bursts associated with binary neutron star mergers are expected to occur at almost the exact time of the merger (cf. Abbott:2017ApJGW170817GRB170817A; Abbott:2017ApJGW170817Multimessenger), providing tight constraints on the time period of data to be searched (of order seconds) for gravitational waves (e.g., Abbott:2021ApJLVKGRBO3a; Abbott:2022ApJLVKGRBO3b). However, for a kilonova candidate detected by Rubin/Fink, the first detection will likely not correspond to the time of the associated neutron star merger. In the worst case scenario, a detection after peak brightness when the transient has already begun to fade, we know the peak must have been some time between the last non-detection and the first detection. This provides a window of order the cadence of the survey, which may be up to 5 days. In Section 4.2.3 we showed that most kilonova candidates identified by Fink will typically be detected one to two days after the neutron star merger. Targeted searches for gravitational waves associated with kilonovae will therefore typically need to search days of data, compared to seconds for searches targeting gamma-ray bursts. This leads to a situation that is more similar to targeted searches for gravitational waves from optically detected supernovae, which also typically have on-source windows of several days (e.g., Abbott:2020PRDSNO1O2; Marek:2024PRDSNO3; Abac:2025ApJSN2023ixf). This increases the computational cost of a targeted search, and it is likely that for many events detector data may not exist for the duration that needs to be searched due to the typical duty cycle of $\sim 70$% for gravitational-wave detectors (e.g., Abac:2025GWTC4Intro). We consider the case of GW170817/AT2017gfo as the best case scenario, that is, a kilonova detected before peak brightness, with a well sampled lightcurve, allowing for models to be fit to estimate the time of the merger. AT2017gfo was first detected in the optical approximately 12 hours after the merger time. From the optical data on AT2017gfo alone, the merger time can be estimated approximately (e.g., Villar:2017ApJGW170817CombinedUVOIR).
@@
The other main difference between a gravitational-wave search targeting kilonova candidates and targeted searches for gravitational waves associated with gamma-ray bursts is that our list of targets will suffer from significant contamination. The majority of kilonova candidates identified by our selection criteria are not actually kilonovae associated with binary neutron star mergers, but rather other variables or transients such as supernovae and asteroids, with similar colors and magnitudes to kilonovae in one or two detections. The situation of a targeted search with significant contamination is similar to the search for gravitational waves associated with fast radio bursts, where some significant fraction (up to 100%) of fast radio bursts may originate from events that are not binary neutron star mergers (Abbott:2023ApJLVKFRBO3a).